OpenStudy (anonymous):

Solve the following exponential equation. 6^x-16*6^-x=6 What is the decimal approx. x≈

OpenStudy (mertsj):

Let y = 6^x

OpenStudy (anonymous):

Do I put that into decimal form?

OpenStudy (mertsj):

Put what in decimal form? 6 ?????

OpenStudy (anonymous):

I am supposed to put the equation answer into decimal form

OpenStudy (mertsj):

Do you have an answer?

OpenStudy (anonymous):

no

OpenStudy (mertsj):

Then what do you have that you are wanting to put into decimal form?

OpenStudy (anonymous):

Im trying to figure it out and don't understand

OpenStudy (mertsj):

So I repeat: let y=6^x. Make the substitution.

OpenStudy (anonymous):

Where do you get y?

OpenStudy (mertsj):

It's a substitution I made up to expedite solving the problem.

OpenStudy (anonymous):

Ok, sorry I don't understand....

OpenStudy (mertsj):

Do you see 6^x in your equation?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Erase it and put a y there.

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

The second term is 16(6^-x). Rewrite that. Move the 6^-x to the bottom and change the sign of the exponent.

OpenStudy (anonymous):

What do you mean move it to the bottom?

OpenStudy (anonymous):

To the end of the equation?

OpenStudy (mertsj):

\[a ^{-n}=\frac{1}{a^n}\]

OpenStudy (mertsj):

Have you ever heard of that property?

OpenStudy (anonymous):

omg I am lost

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I am horrible at this. I have this and another problem I am trying to figure out as well.

OpenStudy (mertsj):

So why do you have to do these problems that are completely beyond your level?

OpenStudy (mertsj):

Have you ever had negative exponents at all?

OpenStudy (anonymous):

All math is beyond me. I have never been good at it. Just not my strong point

OpenStudy (anonymous):

I am pretty sure but I don't remember a lot of it

OpenStudy (mertsj):

\[6^x-\frac{16}{6^x}=6\]

OpenStudy (mertsj):

\[y-\frac{16}{y}=6\]

OpenStudy (mertsj):

Write the left side with a common denominator.

OpenStudy (anonymous):

y=58/7

OpenStudy (mertsj):

I guess I can't figure out where that came from.

OpenStudy (anonymous):

Ok nevermind